I have to problems I need help with
Solve and check
2/5x-1/3x=5
And
If x represents the first three consecutive odd integers, express the sum of the three integers in terms of x

Answer by rfer(16322)

2/5x-1/3x=5
find common denominator
6/15x-5/15x=5
1/15x=5
invert and multiply
x=5/1*15/1
x=75/1
x=75